In 2011, skateboarding legend Micheal Jordan's age was twice that of Olympic Star Lebron James minus 9 years. The sum of their ages in 2011 was 72. How old were Micheal and Lebron in 2011?
Answer by oberobic(2304)

M = Michael Jordan's age in 2011
L = Lebron James' age in 2011
.
M = 2L -9
.
M+L = 72
so
M = 72-L
.
substitute
.
72-L = 2L -9
-3L = -81
L = 27
.
M = 72-27
M = 45
.
Always check your work.
Is M = 2*L -9?
2L = 2*27 =54
54 -9 =45
Yes.
.
Answer: In 2011 Michael was 45 years old, and Lebron was 27 years old.
.
Done.